OpenStudy (anonymous):

Which is the best estimate for the product of 103.132 × 1,031.2? A. 1,000,000 B. 100,000 C. 10,000 D. 1,000

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

its between B D

OpenStudy (skullpatrol):

Why?

OpenStudy (mathmale):

@countsrnc : Kindly explain your reasoning. Do not give out direct answers, even if you're only narrowing down the choices to two from four.

OpenStudy (mathmale):

Better yet, involve @countsrnc in finding his or her own estimate.

OpenStudy (skullpatrol):

Hint: what does 100 X 1, 000 = ?

OpenStudy (anonymous):

100000 = to B 10000

OpenStudy (mathmale):

What have you learned from doing this? What will you do next time you see a problem like this one?

OpenStudy (mathmale):

What value does skullpatrol's suggestion have here? (It's a good one.)

OpenStudy (anonymous):

count the 0s after decimal point and drop one for the answer

OpenStudy (skullpatrol):

$$\huge 100 = 10^2$$ $$\huge 1,000 =10^3$$ $$\huge 100, 000 = 10^{2+3} = 10^5$$